OpenStudy (anonymous):

sin^3theta=-1/8 (find the value of theta for 0

OpenStudy (anonymous):

Let's theta = 0 as symbol [ Sin(0) = -1/2 ]^3 [ 0 =Arcsin (-1/2) ]^3 [ 0 = -30 dgree ]^3 0^3 = - 90 dgree i guess it's